OpenStudy (anonymous):

What is the total cost to repay a credit card loan with a balance of $25,000, an APR of 15%, and a minimum monthly payment of $330?

OpenStudy (amistre64):

how long does it take to pay it off?

OpenStudy (amistre64):

you got any formula for telling you the current balance of a loan?

OpenStudy (anonymous):

sorry, no

OpenStudy (amistre64):

it would be useful if you had some formula that could tell you such things. Payment formula is to ... ugh Any present value or future value formulas ?

OpenStudy (amistre64):

i have one that i developed during accounting class, the text books were giving all sorts of formulas and so i just developed my own that pretty much combined them all

OpenStudy (anonymous):

what is it?

OpenStudy (amistre64):

spose you have balance, B, compounding at some rate, k, over n periods Bn = Bk^n now spose the balance changes due to some regular payment, P. The sum of the payments made follow a geometric series whose sum is the formula: (k^n-1)/(k-1) therefore, lets subtract our payments as Bn = Bk^n - P(k^n-1)/(k-1) when Bn = 0, the loan is paid right?

OpenStudy (anonymous):

...I think so...

OpenStudy (amistre64):

do we like paying loans after they get to a 0 balance? of course not :) Bn=0 is the loan paid in full so we just solve for n to know how long it takes

OpenStudy (amistre64):

Bn = Bk^n - P(k^n-1)/(k-1) 0 = Bk^n - P(k^n-1)/(k-1) 0 = Bk^n(k-1) - P(k^n-1) 0 = Bk k^n - Bk^n - Pk^n -1 0 = k^n(Bk - B - P) -1 1/(Bk - B - P) = k^n log(1/(Bk - B - P)) = n log(k) - log((Bk - B - P))/log(k) = n

OpenStudy (amistre64):

the total cost of the loan is n*P

OpenStudy (anonymous):

...so confused

OpenStudy (amistre64):

im not a mind reader so youll have to express that confusion in a manner that i can actually aid in.

OpenStudy (amistre64):

do we have a formula developed to find the total number of payments needed?

OpenStudy (anonymous):

what the heck are 'b' and 'p' and 'k' and 'n'?????

OpenStudy (amistre64):

i defined them clearly in the postings balance, payments, number of periods, an compounding interest

OpenStudy (amistre64):

as previously posted above: _____________________________ spose you have balance, B, compounding at some rate, k, over n periods Bn = Bk^n now spose the balance changes due to some regular payment, P. The sum of the payments made follow a geometric series whose sum is the formula: (k^n-1)/(k-1)